# Density-based clustering algorithm for galaxy group/cluster identification

### v1

@@ 5.1 Performance of finding BCGs
In employing the scaled OPTICS clustering algorithm to identify galaxy clusters, the hyperparameters $\epsilon$ (the maximum radius for neighborhood density estimation), $N_{\mathrm{min}}$ (the minimum number of points required to form a cluster), and the LOS scaling factor crucially influence the results, as detailed in Section 3. These parameters are pivotal in determining reachability distances and adjusting the algorithm to mitigate the effects of redshift-space distortion.
Although one might expect the clustering outcome to be highly sensitive to parameter variations, sOPTICS shows resilience by exhibiting an optimal range for their values. This finding is illustrated in Figures 9 and 10, where the interdependence of $\epsilon$ and $N_{\mathrm{min}}$, as well as $\epsilon$ and $s_{\mathrm{LOS}}$, is presented. Notably, a clear correlation emerges between $\epsilon$ and $N_{\mathrm{min}}$; as $\epsilon$ increases, $N_{\mathrm{min}}$ also needs to be adjusted upward to maintain effective clustering. Essentially, to preserve the purity of the clusters identified by the sOPTICS algorithm, the criteria must shift toward identifying denser and larger clusters as the $\epsilon$ threshold is raised. The positive correlation suggests that $\epsilon^{3}\propto N_{\mathrm{min}}$ aligns with theoretical expectations. Theoretically, this adjustment ensures that the increase in the neighborhood radius does not lead to the inclusion of outlier points or less dense areas. Thus $N_{\mathrm{min}}$ should correlate with the volume of space encompassed within $\epsilon$, which implies a cubic relationship ($\epsilon^{3}$).
Figure 9: The results of clustering for various pairs of values of $\epsilon$ and $N_{\mathrm{min}}$. The color bar represents the weighted recovery rate of the predicted groups relative to a subsample from the Y07 groups catalog, covering a $30^{2}\,\deg^{2}$ sky area within the redshift range $0.80.7<z<$ 0.8.
Figure 10: As Fig. 9, but showing the results of clustering for various pairs of values of $s_{\mathrm{LOS}}$ and $\epsilon$.
These relationships underscore the interconnected nature of these parameters and their collective impact on optimizing cluster detection and recovery rates. They are fundamental in understanding how changes in one parameter necessitate adjustments in the other to maintain the purity and completeness of clustering outcomes. Specifically, within a certain redshift range, selecting values for these three parameters from this optimal range can yield galaxy groups with a high recovery rate comparable to those obtained from reliable group catalogs that require complex and computationally intensive processes. The optimal ranges depicted in Figures 9 and 10 also identify a potential characteristic number density for categorizing galaxies in a survey as a group. Given the similar local completeness of the survey, this characteristic number density of galaxy groups can be applied to other observations. This consideration is pivotal when choosing hyperparameter values.
However, given the current redshift range of the observed data, the relationships observed between $\epsilon$, $N_{\mathrm{min}}$, and $s_{\mathrm{LOS}}$ are preliminary and roughly empirical. To more precisely define these relationships and understand the characteristic number density, a comprehensive analysis using both real-world data and mock catalogs is crucial. This approach would help determine whether the observed linear trend between $\epsilon$ and $s_{\mathrm{LOS}}$ is an artifact of the specific dataset used in this study or if it reflects a more general characteristic applicable across different galaxy cluster distributions. Furthermore, a theoretical exploration into the dynamics of galaxy clusters and their spatial density distributions, similar to the approach taken by Y07, should be considered. Such studies would provide deeper insights into the viability of sOPTICS methods and enrich our understanding of the underlying physical properties that govern the formation and evolution of galaxy clusters.
In addition, given that BCGs are typically located near the densest part of a cluster, sOPTICS, which specifically focuses on dense patterns within a survey, naturally excels as a tool for locating BCGs. It indeed has shown promising results in tests against the Hsu22 sample (see Section 5.1. Therefore, for researchers primarily interested in identifying BCGs, sOPTICS can be effectively employed with a specific set of parameters tailored for very dense groups: small values of $\epsilon$, large $N_{\mathrm{min}}$, and reasonable values of $s_{\mathrm{LOS}}$. This setup allows for rapidly identifying potential BCG candidates from large and complex surveys. In contrast, the FoF can not efficiently identify purely large clusters without contaminating smaller ones due to small values for the linking length. Still, It is important to note that while BCGs are the most luminous galaxies within their clusters, they may not always coincide with the cluster’s geometric center, which is influenced by both the geometry and the luminosity of member galaxies (Skibba et al., 2011). Accurate identification of BCGs requires a meticulous selection process that considers factors such as brightness, proximity to the cluster center, distinctive features, and corroborative data from various observational sources. Subsequent verification using additional observations, such as X-ray and optical surveys, should be done on the candidates to confirm the true BCGs.
@@ 6 Summary and Conclusion
This study evaluated the effectiveness of eight popular clustering algorithms in data science for identifying galaxy groups and clusters through tests involving comparisons with both simulations and existing reliable group catalogs. Our findings indicate that the FoF and sOPTICS algorithms are robust galaxy group finders. Remarkably, our sOPTICS method exhibits considerable stability and flexibility in the choice of hyperparameter values and, enhanced by the line-of-sight scaling factor to mitigate redshift-space distortion, outperforms FoF in both efficiency and accuracy in identifying the most geometrically dense parts of galaxy groups and in pinpointing BCGs from large surveys.
We conclude that scaled OPTICS and FoF are comparably effective, with sOPTICS demonstrating high purity and recovery rates. While FoF can be faster and more computationally efficient, especially for large datasets – an advantage in astronomical computations – its performance heavily depends on the choice of linking length. Despite this dependency, FoF, as a popular and classical clustering method in astrophysics, remains particularly effective for low redshift surveys where redshift space distortion is less significant.
sOPTICS is particularly noteworthy for three reasons: i) It demonstrates robustness to a wide range of hyperparameter values. Most importantly, there are two empirical relationships involving its hyperparameters $\epsilon$, $N_{\mathrm{min}}$, and $s_{\mathrm{LOS}}$ enabling us to estimate a reasonable range of values to achieve reliable clustering results. ii) Unlike many clustering algorithms, sOPTICS does not segment data into clusters during its initial run. Instead, it generates a reachability plot illustrating distances to the nearest neighbor within the $\epsilon$-neighborhood. Clusters are subsequently identified based on the valleys within this plot. This approach allows sOPTICS to be less sensitive to hyperparameter settings as long as they are sufficient to capture significant structures in the data. iii) sOPTICS primarily focuses on the densest parts of a dataset, where the BCGs are often located. This focus makes it naturally efficient in identifying BCGs. By setting even extreme hyperparameter values, researchers can easily identify reliable BCG candidates for further analysis, simplifying the process compared to more complex galaxy group-finding methods.
Looking ahead, we anticipate leveraging richer and more precise galaxy data from sources such as the Dark Energy Spectroscopic Instrument (DESI, Dey et al., 2019) and other ongoing large-scale spectroscopic surveys. This will enable more realistic treatments in hyperparameter modeling, especially for galaxy groups and clusters at higher redshifts. Furthermore, we aim to refine the relationships and potential characteristic numbers suggested by the hyperparameters $\epsilon$, $N_{\mathrm{min}}$, and $s_{\mathrm{LOS}}$ in sOPTICS.
@@ Acknowledgements
The data underlying this paper is publicly available at http://sdss.physics.nyu.edu/vagc/ and https://gax.sjtu.edu.cn/data/Group.html. The derived data generated in this research will be shared on a reasonable request to the first author.

### v2

@@ 5.3 Performance of finding BCGs
One of the key strengths of sOPTICS is its ability to serve as a proxy for those complex physically motivated, FoF-based approaches while retaining efficiency and adaptability. By focusing on dense regions in a dataset, sOPTICS identifies galaxy clusters across varying environments with minimal tuning and computational overhead, as demonstrated by its high recovery rate relative to the Y07 catalog and other observational samples (Sections 5.2). These findings highlight sOPTICS’s utility in constructing reliable galaxy group catalogs for large redshift surveys. Even when parameters are not extensively recalibrated, sOPTICS can maintain strong performance, offering both practical scalability and solid clustering outcomes. This resilience across diverse datasets makes sOPTICS a promising, and efficient tool for future galaxy-group identification tasks, especially in scenarios where redshift-space distortions pose significant challenges. Unlike FoF, which requires carefully chosen linking lengths that can merge or dilute cluster boundaries, sOPTICS adaptively captures complex group structures and provides valuable information on internal density variations within clusters.
An additional advantage of sOPTICS lies in its natural compatibility with BCGs searches. Because BCGs generally reside near the densest region of a cluster, a clustering algorithm that prioritizes dense structures can be particularly effective at locating these galaxies. Our tests against the Hsu22 sample confirm this potential (Section 5.3). For researchers specifically targeting BCGs, we recommend using sOPTICS with relatively small $\epsilon$, large $N_{\mathrm{min}}$, and reasonable $s_{\mathrm{LOS}}$ values to highlight very dense clusters. This setup streamlines the search for potential BCG candidates in expansive surveys, offering a time-efficient alternative to highly iterative or computationally intensive methods.
Nevertheless, while BCGs are often the most luminous galaxies within a cluster, they do not necessarily coincide with the cluster’s geometric center, given that spatial centering is influenced by halo geometry and luminosity distributions (Skibba et al., 2011). Hence, accurately identifying a BCG requires a careful selection process that considers factors such as brightness, proximity to the cluster center, and corroborating information from multi-wavelength surveys. Verification with X-ray or optical follow-up can further constrain the reliability of a given BCG identification.
Taken together, our findings demonstrate that sOPTICS can be used effectively on both simulated and real datasets to identify large-scale structures and delve into their internal density patterns. Its efficiency, adaptability, and capacity for uncovering the densest regions within clusters make it a promising choice for both general group-finding purposes and specialized tasks such as BCG identification.
@@ 6 Summary and Conclusion
This study evaluated the effectiveness of eight popular clustering algorithms in data science for identifying galaxy groups and clusters through tests involving comparisons with both simulations and existing reliable group catalogs. Our findings indicate that our sOPTICS algorithm is a robust galaxy group finders. In particular, sOPTICS demonstrates significant flexibility in its hyperparameters and, when combined with a line-of-sight scaling factor to mitigate redshift-space distortion, exceeds FoF in both efficiency and accuracy. This advantage is especially apparent in pinpointing the densest regions of galaxy groups and identifying BCGs across large surveys.
We conclude that scaled OPTICS and FoF are comparably effective, with sOPTICS showing higher purity and recovery rates. While FoF can be faster and more computationally efficient, especially for large datasets – a recognized asset for very large datasets from observation – its performance heavily depends on the choice of linking length. Despite this dependency, FoF, as a popular and classical clustering method in astrophysics, remains particularly effective for low redshift surveys where redshift space distortion is less significant.
Our investigation highlights three principal strengths of sOPTICS:
  • Robustness to a wide range of hyperparameter values. We have identified two empirical relationships involving $\epsilon$, $N_{\mathrm{min}}$, and $s_{\mathrm{LOS}}$, which provide practical guidance for setting these parameters to achieve reliable clustering results.
[…]
  • Focusing on the densest regions of the dataset, where BCGs frequently reside. By choosing extreme but physically motivated hyperparameter values, one can efficiently isolate the most significant clusters, making sOPTICS a practical method for surveying massive structures without requiring more complex group-finding computing.
[…]
@@ Acknowledgements
The data utilized in this study is publicly accessible at the following sources: http://sdss.physics.nyu.edu/vagc/ and https://gax.sjtu.edu.cn/data/Group.html. Additionally, the code and derived data produced during this research can be made available upon reasonable request to the first author.